calculate the pH of a 0.0105M solution of NaOH?
1
Expert's answer
2012-05-08T12:01:26-0400
pH=-14-pOH
pOH=-log[OH-]
pH=14-(-log[OH-])=14-(-log[0.0105])=14-1,98=12.02